write the equation of a line that passes through (1,-7) with a slope of 2/3. I need help solving this problem and the steps on how to solve it so I can try to figure out the other ones just like it! thanks!
Answer by mananth(16946)

m= 2/ 3

Plug value of the slope and point (1 ,-7) in to find value of b
Y = m x + b
-7.00 = 2/3 + b
b= -7 - 2/3
b= -23/ 3
So the equation will be
Y = 2/3 x -23/3
m.ananth@hotmail.ca